# Frequency-Guided Posterior Sampling for Diffusion-Based Image Restoration

### v1

@@ Appendix C Further Motivations for our Method
Appendix D Experimental Details
@@
D.1 Inverse Problem Task Descriptions
D.1.1 Linear Inverse Problems
@@
D.1.2 Nonlinear Inverse Problems
@@ D.2 Baselines
Score-SDE/ILVR [35, 8] Generally, we group Score-SDE and ILVR as methods that use a sequence of noisy measurements to approximate the conditional score, as mentioned in [12] and [9]. This is a generalization of the methods in [35] and [8] as their methods only were presented for the inpainting and superresolution tasks. To consider general tasks, we sample a sequence of measurements $\mathbf{y}_{t}\sim\mathcal{N}(\sqrt{\bar{\alpha_{t}}}\mathbf{y},(1-\bar{\alpha_{t}}\mathbf{I}))$. Then, we approximate the conditional score as $-\eta\nabla_{\mathbf{x}_{t}}\left\|\mathbf{y}_{t}-\mathcal{A}(\mathbf{x}_{t})\right\|_{2}^{2}$. We use the step size $\eta$ from the DPS method [9].
@@
D.3 Details for Figure 2 and Figure 3
In Figure 2 and Figure 3, we demonstrated the approximation gap of DPS and our method on synthetic data. Below, we describe the precise experimental details of our experiments.
Forward Operators. We consider two operators in our experiments: a low-pass filter and a high-pass filter. For the low-pass filter, we create a circulant matrix $\mathbf{A}_{l}$ whose first row is a Gaussian blur kernel of width $\sigma$ i.e. this kernel is $e^{-\mathbf{x}^{2}/(2\sigma^{2})}$ and normalized to have sum $1$. Multiplying $\mathbf{A}_{l}$ with a signal $\mathbf{x}$ corresponds to convolving $\mathbf{x}$ with the Gaussian blur kernel. In our experiments, we vary $\sigma$ as shown as the different lines in the figures. For the high-pass filter, we simply subtract the Gaussian blur kernel of width $\sigma$ from a Dirac kernel, which has 1 at the center and zeroes elsewhere. Importantly, the first row of the circulant matrix corresponding to this kernel is normalized to have sum zero, such that the kernel is indeed a high-pass filter.
@@ D.3 Details for Figure 2 and Figure 3
Approximation Gap. For each $\mathbf{x}^{(i)},\mathbf{y}^{(i)}$ pair, we calculate $\mathbf{x}_{t}^{(i)}\sim\mathcal{N}(\sqrt{\bar{\alpha_{t}}}\mathbf{x}^{(i)},(1-\bar{\alpha_{t}})\mathbf{I})$, where $\bar{\alpha_{t}}$ is the variance schedule from the DDPM paper [20]. We report the average approximation gap over the 10000 signals.
Frequency Curriculum. For our method, we have the choices of the frequency curriculum described through $\tau_{T}$ and $\tau_{1}$, the frequency cutoffs for the time-dependent low pass filter. For simplicity, we use a frequency curriculum from the variance schedule of the diffusion model. Specifically, let $\sigma_{t}=\frac{1}{\bar{\alpha_{t}}}-1$, which corresponds to the corresponding variance-exploding parameterization of the SDE (see [24], Appendix B). Then we set $\tau_{t}$ be the max frequency value $f_{k}$ such that $S(f_{k})\geq\max\{\sigma_{y}^{2},\sigma_{t}^{2}\}$, where $S(f_{k})$ denotes the power spectral density from Equation 8. This utilizes the intuition that for $\mathbf{x}_{t}^{(i)}$, the frequency components of the original signal $\mathbf{x}^{(i)}$ still present in the image are the frequencies below $\tau_{t}$.
D.4 Implementation Details for Our Method
@@
In Appendix F, we provide ablation studies to understand the effects of these hyperparameters on the generation quality.
D.5 Computational Overhead of Our Method
In this section, we evaluate the computational efficiency of our method by measuring the average runtime per image on the FFHQ dataset. We run each method from Table 1 on 100 images, and report the average time taken per image, given in Table 3. We see that our method introduces minimal computational overhead compared to the DPS method, demonstrating that our modifications can be implemented in an efficient manner and easily integrated into existing frameworks.
@@ D.5 Computational Overhead of Our Method
Appendix E Comparison to ILVR/Score-SDE
Our method is closely related to the Score-SDE and ILVR works that consider a sequence of noisy measurements $y_{t}$ such that $y_{t}=\mathcal{N}(\sqrt{\alpha_{t}}y_{0},(1-\alpha)\mathbf{I})$ [8, 35]. These methods consider approximations to the noisy likelihood score that are typically of the form $\mathbf{L}_{\mathcal{A}}(\mathbf{y}_{t}-\mathcal{A}(\mathbf{x}_{t}))$, where $\mathbf{L}_{\mathcal{A}}$ is a fixed matrix that depends on the measurement operator $\mathcal{A}$. There are three crucial differences between these methods and our method. First, our approximation to the noisy likelihood score considers a time-varying model likelihood at each diffusion timestep $p(\mathbf{y}\mid\bm{\mu}_{0\mid t})$ as given in Equation 13 as opposed to simply varying $\mathbf{y}_{t}$. Further, we leverage the powerful denoising capabilities of the pretrained diffusion model by using the Tweedie estimate $\bm{\mu}_{0\mid t}$ instead of the noisy $\mathbf{x}_{t}$. Lastly, we propose a frequency curriculum that can differ from the variance schedule of the diffusion model and be adapted to the frequency characteristics of the data. These differences lead to a drastically improved performance of our method compared to Score-SDE/ILVR as seen in Table 1.
Appendix F Further Experiments and Ablation Studies
F.1 Visualizing the Transformed Measurements
In Figure 7, we show the transformed measurements at three timesteps in the reverse diffusion process when applying our frequency curriculum on the FFHQ dataset. As we can see, in the early stages of the reverse process, the measurement retains only a coarse outline of the measurement. Our theoretical results indicate that the initial reverse process steps are very important in order to obtain coarse alignment with the given measurement.
@@ F.1 Visualizing the Transformed Measurements
F.2 Effect of Frequency Curriculum
Impact of Time-Dependent Curriculum. As observed in Figure 5, we observe that the time-dependent frequency curriculum helps the stability of the method. Namely, when the operator is a high pass filter, the time-dependent frequency curriculum results in high-quality reconstructions. On the other hand, for a fixed curriculum, we observe that over different generations, most random samples of the diffusion model fall off the natural image manifold in the early steps of the reverse process, which results in unnatural looking images as in Figure 5.
@@ F.2 Effect of Frequency Curriculum
F.3 Effect of Step Size Schedule
Recall that in Equation 15, we had that our approximation to the score should be scaled by $\mathbf{S}_{t}$, which can be thought of as a step size to the gradient terrm. In practice, instead of using this exact quantity, which would require using a potentially large circulant matrix and its inverse, we set $\mathbf{S}_{t}=\frac{\kappa}{\left\|\phi_{t}(\mathbf{y})-\phi(\mathcal{A}(\bm{\mu}_{0\mid t}))\right\|_{2}}\mathbf{I}$ for a scalar hyperparameter $\kappa$, which works well in practice. This is similar to the step size chosen in DPS. The intuition behind this step size choice is to control the approximation error. For example, when the approximation error is high, the step size should be smaller as the conditional score is noisy. These schedules play a crucial role in the effectiveness and stability of the optimization process. In this section, we study the effect of different step size schedules to motivate our choice. Specifically, we study three step size schedules (where we vary $\kappa$) on the motion deblurring task on the FFHQ dataset. These three schedules are:
@@ F.3 Effect of Step Size Schedule
    Cosine Annealed step size according to Equation 50. This schedule starts with a relatively high step size, which facilitates rapid progress in the initial stages, and gradually decreases to smaller step sizes as the optimization nears convergence.
In Table 4 and Figure 8, we demonstrate the quantitative and qualitative differences between the three schedules. We clearly see that the cosine annealed step size strongly outperforms the other two schedules, which is why we adopt it for our experiments. We hypothesize that the cosine schedule allows the model to quickly capture the coarse structure of the image early on, while the progressively smaller steps enable fine-tuning of details, leading to a more refined final reconstruction. This is evident from Figure 8. For example, in the first row, the cosine schedule successfully captures the necklace detail around the subject’s neck, while the fixed small step size schedule completely misses this feature, resulting in a blurred and oversmoothed output. The fixed large step size schedule manages to retain some necklace details but introduces noticeable artifacts, which degrade the overall image quality. The large step size likely destabilizes the optimization and pushes the image off the natural image manifold, which compromises image quality.
@@
F.4 Further Qualitative Results
Comparison to Baselines. In Figure 11, we see the zoomed in version from Figure 4, highlighting the intricate details of the image captured by our reconstructions. In Figure 12, we see a comprehensive comparison to all the baselines reported in Table 1 on the motion blurring task on the FFHQ dataset. We observe that Score-SDE/ILVR often captures the correct structure of the image but fills in different facial features. MCG usually overfits to the Gaussian measurement noise as reported in [9]. DPS results in good quality reconstructions, but the images are usually smoothed out and lose small facial features. DSG performs very well on the motion deblurring task, as shown also in Table 1. However, DSG often gives grainy reconstructions and artifacts that are clearly visible. In contrast, our method is significantly more stable and is able to give images that have higher perceptual quality than all the baselines.
@@ F.4 Further Qualitative Results
ImageNet Additional Results. Figures 15, 16, and 17 show additional results of our method on the Gaussian deblurring, motion deblurring, and high-pass filter deconvolution tasks on the ImageNet dataset. On the high-pass filter task, we note that there is sometimes a color shift which results from the loss of color information in the measurement. We observe that compared to DPS and a fixed-time frequency curriculum (as in Figure 5), our method is much more stable and usually gives visually plausible reconstructions instead of strong color artifacts that dominate the reconstruction.
Figure 11: Qualitative results of our method with zoomed in portions of images from Figure 4. Our method successfully preserves finer details like background pattern.
Figure 12: Qualitative motion deblurring results on FFHQ dataset for all baselines we report in Table 1. The same blur kernel is applied to each image.
@@
Appendix G Limitations and Future Work

### v2

@@ Appendix D Further Motivations for our Method
Lastly, we emphasize that our experimental results demonstrate a fascinating phenomenon where the performance gap between FGPS and baseline methods is even larger on more complex datasets, hinting at the fact that besides the frequency characteristics of the forward operator, the frequency characteristics and quality of the Tweedie estimate also greatly affects reconstruction. It would be an interesting theoretical direction to understand why the frequency schedule has larger benefits in this case.
Appendix E Experimental Details
@@
E.1 Inverse Problem Task Descriptions
E.1.1 Linear Inverse Problems
@@
E.1.2 Nonlinear Inverse Problems
@@ E.2 Baselines
Score-SDE/ILVR [37, 8] Generally, we group Score-SDE and ILVR as methods that use a sequence of noisy measurements to approximate the conditional score, as mentioned in [12] and [9]. This is a generalization of the methods in [37] and [8] as their methods only were presented for the inpainting and superresolution tasks. To consider general tasks, we sample a sequence of measurements $\mathbf{y}_{t}\sim\mathcal{N}(\sqrt{\bar{\alpha_{t}}}\mathbf{y},(1-\bar{\alpha_{t}}\mathbf{I}))$. Then, we approximate the conditional score as $\eta\nabla_{\mathbf{x}_{t}}\left\|\mathbf{y}_{t}-\mathcal{A}(\mathbf{x}_{t})\right\|_{2}^{2}$. We use the step size $\eta$ from the DPS method [9].
@@
E.3 Details for Figure˜1
In Figure˜1, we demonstrated the approximation gap of DPS and our method on synthetic data. Below, we describe the precise experimental details of our experiments.
Forward Operators. We consider a high-pass filter operators in our experiments. a low-pass filter and a high-pass filter. First, we create a Gaussian blur kernel of width $\sigma$ i.e. this kernel is $e^{-\mathbf{x}^{2}/(2\sigma^{2})}$ and normalized to have sum $1$. For the high-pass filter, we simply subtract the Gaussian blur kernel of width $\sigma$ from a Dirac kernel, which has 1 at the center and zeroes elsewhere. The filter is a circulant matrix constructed from this kernel. Importantly, the first row of the circulant matrix corresponding to this kernel is normalized to have sum zero, such that the kernel is indeed a high-pass filter.
@@ E.3 Details for Figure˜1
Approximation Gap. For each $\mathbf{x}^{(i)},\mathbf{y}^{(i)}$ pair, we calculate $\mathbf{x}_{t}^{(i)}\sim\mathcal{N}(\sqrt{\bar{\alpha_{t}}}\mathbf{x}^{(i)},(1-\bar{\alpha_{t}})\mathbf{I})$, where $\bar{\alpha_{t}}$ is the variance schedule from the DDPM paper [20]. We report the average approximation gap over the 10000 signals. Below, we give the exact approximation that our method, FGPS, makes for the noisy likelihood score. From Theorem˜2, we know that under a multivariate normal assumption on the data, the posterior mean $\bm{\mu}_{0\mid t}$ is $\bm{\Gamma}_{t}\mathbf{x}_{t}$. Under the same assumptions as Theorem˜2, for linear inverse problems where $\mathcal{A}(\mathbf{x}_{0})=\mathbf{A}\mathbf{x}_{0}$ we analytically compute our approximation to the noisy likelihood score given in Equation˜12 as
Frequency Curriculum. For our method, we have the choices of the frequency curriculum described through $\tau_{T}$ and $\tau_{1}$, the frequency cutoffs for the time-dependent low pass filter. For simplicity, we use a frequency curriculum from the variance schedule of the diffusion model. Specifically, let $\sigma_{t}=\frac{1}{\bar{\alpha_{t}}}-1$, which corresponds to the corresponding variance-exploding parameterization of the SDE (see [24], Appendix B). Then we set $\tau_{t}$ be the max frequency value $f_{k}$ such that $S(f_{k})\geq\max\{\sigma_{y}^{2},\sigma_{t}^{2}\}$, where $S(f_{k})$ denotes the power spectral density from Equation˜8. This utilizes the intuition that for $\mathbf{x}_{t}^{(i)}$, the frequency components of the original signal $\mathbf{x}^{(i)}$ still present in the image are the frequencies below $\tau_{t}$.
E.4 Implementation Details for Our Method
@@
In Appendix G, we provide ablation studies to understand the effects of these hyperparameters on the generation quality.
E.5 Computational Overhead of Our Method
In this section, we evaluate the computational efficiency of our method by measuring the average runtime per image on the FFHQ dataset. We run each method from Table˜1 on 100 images, and report the average time taken per image, given in Table 3. We see that our method introduces minimal computational overhead compared to the DPS method, demonstrating that our modifications can be implemented in an efficient manner and easily integrated into existing frameworks.
@@ E.5 Computational Overhead of Our Method
Appendix F Comparison to ILVR/Score-SDE
Our method is closely related to the Score-SDE and ILVR works that consider a sequence of noisy measurements $y_{t}$ such that $y_{t}=\mathcal{N}(\sqrt{\alpha_{t}}y_{0},(1-\alpha)\mathbf{I})$ [8, 37]. These methods consider approximations to the noisy likelihood score that are typically of the form $\mathbf{L}_{\mathcal{A}}(\mathbf{y}_{t}-\mathcal{A}(\mathbf{x}_{t}))$, where $\mathbf{L}_{\mathcal{A}}$ is a fixed matrix that depends on the measurement operator $\mathcal{A}$. There are three crucial differences between these methods and our method. First, our approximation to the noisy likelihood score considers a time-varying model likelihood at each diffusion timestep $p(\mathbf{y}\mid\bm{\mu}_{0\mid t})$ as given in Equation˜10 as opposed to simply varying $\mathbf{y}_{t}$. Further, we leverage the powerful denoising capabilities of the pretrained diffusion model by using the Tweedie estimate $\bm{\mu}_{0\mid t}$ instead of the noisy $\mathbf{x}_{t}$. Lastly, we propose a frequency curriculum that can differ from the variance schedule of the diffusion model and be adapted to the frequency characteristics of the data. These differences lead to a drastically improved performance of our method compared to Score-SDE/ILVR as seen in Table˜1.
Appendix G Further Experiments and Ablation Studies
G.1 Visualizing the Transformed Measurements
In Figure˜7, we show the transformed measurements at three timesteps in the reverse diffusion process when applying our frequency curriculum on the FFHQ dataset on all the forward operators we considered. Our theoretical results indicate that the initial reverse process steps are very important in order to obtain coarse alignment with the given measurement, so it is reasonable to align the measurements with this coarse-to-fine strategy.
@@ G.1 Visualizing the Transformed Measurements
G.2 Effect of Frequency Curriculum
Impact of Time-Dependent Curriculum. As observed in Figure˜5, we observe that the time-dependent frequency curriculum helps the stability of the method. Namely, when the operator is a high pass filter, the time-dependent frequency curriculum results in high-quality reconstructions. On the other hand, for a fixed curriculum, we observe that over different generations, most random samples of the diffusion model fall off the natural image manifold in the early steps of the reverse process, which results in unnatural looking images as in Figure˜5.
@@ G.2 Effect of Frequency Curriculum
G.3 Effect of Step Size Schedule
Recall that in Equation˜12, we had that our approximation to the score should be scaled by $\mathbf{S}_{t}$, which can be thought of as a step size to the gradient terrm. In practice, instead of using this exact quantity, which would require using a potentially large circulant matrix and its inverse, we set $\mathbf{S}_{t}=\frac{\kappa}{\left\|\phi_{t}(\mathbf{y})-\phi(\mathcal{A}(\bm{\mu}_{0\mid t}))\right\|_{2}}\mathbf{I}$ for a scalar hyperparameter $\kappa$, which works well in practice. This is similar to the step size chosen in DPS. The intuition behind this step size choice is to control the approximation error. For example, when the approximation error is high, the step size should be smaller as the conditional score is noisy. These schedules play a crucial role in the effectiveness and stability of the optimization process. In this section, we study the effect of different step size schedules to motivate our choice. Specifically, we study three step size schedules (where we vary $\kappa$) on the motion deblurring task on the FFHQ dataset. These three schedules are:
@@ G.3 Effect of Step Size Schedule
    Cosine Annealed step size according to Equation˜48. This schedule starts with a relatively high step size, which facilitates rapid progress in the initial stages, and gradually decreases to smaller step sizes as the optimization nears convergence.
In Table˜4 and Figure˜8, we demonstrate the quantitative and qualitative differences between the three schedules. We clearly see that the cosine annealed step size strongly outperforms the other two schedules, which is why we adopt it for our experiments. We hypothesize that the cosine schedule allows the model to quickly capture the coarse structure of the image early on, while the progressively smaller steps enable fine-tuning of details, leading to a more refined final reconstruction. This is evident from Figure˜8. For example, in the first row, the cosine schedule successfully captures the necklace detail around the subject’s neck, while the fixed small step size schedule completely misses this feature, resulting in a blurred and oversmoothed output. The fixed large step size schedule manages to retain some necklace details but introduces noticeable artifacts, which degrade the overall image quality. The large step size likely destabilizes the optimization and pushes the image off the natural image manifold, which compromises image quality.
@@
G.4 Further Qualitative Results
Comparison to Baselines. In Figure˜11, we see the zoomed in version from Figure˜3, highlighting the intricate details of the image captured by our reconstructions. In Figure˜12, we see a comprehensive comparison to all the baselines reported in Table˜1 on the motion blurring task on the FFHQ dataset. We observe that Score-SDE/ILVR often captures the correct structure of the image but fills in different facial features. MCG usually overfits to the Gaussian measurement noise as reported in [9]. DPS results in good quality reconstructions, but the images are usually smoothed out and lose small facial features. DSG performs very well on the motion deblurring task, as shown also in Table˜1. However, DSG often gives grainy reconstructions and artifacts that are clearly visible. In contrast, our method is significantly more stable and is able to give images that have higher perceptual quality than all the baselines.
@@ G.4 Further Qualitative Results
ImageNet Additional Results. Figures 15, 16, and 17 show additional results of our method on the Gaussian deblurring, motion deblurring, and high-pass filter deconvolution tasks on the ImageNet dataset. On the high-pass filter task, we note that there is sometimes a color shift which results from the loss of color information in the measurement. We observe that compared to DPS and a fixed-time frequency curriculum (as in Figure˜5), our method is much more stable and usually gives visually plausible reconstructions instead of strong color artifacts that dominate the reconstruction.
Figure 11: Qualitative results of our method with zoomed in portions of images from Figure˜3. Our method successfully preserves finer details like background pattern.
Figure 12: Qualitative motion deblurring results on FFHQ dataset for all baselines we report in Table˜1. The same blur kernel is applied to each image.
@@
Appendix H Limitations and Future Work
@@
It is important to note that our theoretical results do not paint the full picture of the success of FGPS in practice. Empirically, the Tweedie estimate $\bm{\mu}_{0\mid t}$ used for the conditional score approximation behaves in complex ways and its frequency structure is not as simple as the form in our theoretical results, $\bm{\Gamma}_{t}\mathbf{x}_{0}$. We conjecture it is still important to explicitly align the frequency structure of the unconditional score and noisy likelihood score, which is why FGPS outperforms DPS for low-frequency measurements on FFHQ and Imagenet. In addition to the role of the Tweedie estimate, we find an intriguing role of the dataset where FGPS performs even better on harder datasets like ImageNet due to more complex frequency structure. Explaining both these phenomena theoretically is an interesting direction for future work. That being said, our empirical findings demonstrate that the core idea behind FGPS, aligning the spectral structure of the measurement with the score function, remains effective for complex data. Our curriculum strategy reflects a coarse-to-fine alignment of frequencies, motivated by both the empirical behavior of diffusion models and spectral properties of natural images. We believe this insight opens avenues for more principled guidance mechanisms utilizing the structure of the score function.